The population standard deviation for the height of high school basketball players is 3.3 inches. If we want to be 95% confident that the sample mean height is within 1.8 inch of the true population mean height, how many randomly selected students must be surveyed? Fill in the blank: n=
Answer #1

Answer:

Given that:

The population standard deviation for the height of high school basketball players is 3.3 inches.

population standard deviation=3.3

Margin of error,E=1.8

For 95% confidence,z=1.96

Sample size required,n=(1.96*3.3/1.8)^2

= (6.468/1.8)^2

=(3.5933) ^2

=12.91

n = 12